### 07 Jan 2024 v5

@@ III-E Linear Complexity Intra-Slice Global Context Module
One drawback of vanilla approach is its quadratic complexity. In Equation 16, the softmax operation specifies the order of tensor calculation. The complexity of ${{\hat{\bm{y}}}^{i-1}}_{na,q}\times\left({{\hat{\bm{y}}}^{i-1}}_{a,k}\right)^{\top}$ is $O(L^{2})$. The quadratic complexity leads to huge GPU memory consumption, longer encoding and decoding time as illustrated in Fig. 12, which makes it hard to employ the vanilla approach for high-resolution image coding. In Equation 16, if $\left(\hat{\bm{y}}^{i-1}_{ac,k}\right)^{\top}\times\hat{\bm{y}}^{i}_{ac,v}$ is computed first, the overall complexity becomes $O(L)$, which is linear with the resolution. Equation 16 works because $10<\textrm{softmax}\left(\frac{{{\hat{\bm{y}}}^{i-1}}_{na,q}\times\left({{\hat{\bm{y}}}^{i-1}}_{ac,k}\right)^{\top}}{\sqrt{S}}\right)<$ 1. The non-negativity makes it can be treated as a learnable similarity metric. If $\textrm{softmax}\left(\frac{{{\hat{\bm{y}}}^{i-1}}_{na,q}\times\left({{\hat{\bm{y}}}^{i-1}}_{ac,k}\right)^{\top}}{\sqrt{S}}\right)\to 0$, ${\hat{\bm{y}}}^{i-1}_{na,q}$ and ${\hat{\bm{y}}}^{i-1}_{na,q}$ are near orthogonal. If $\textrm{softmax}\left(\frac{{{\hat{\bm{y}}}^{i-1}}_{na,q}\times\left({{\hat{\bm{y}}}^{i-1}}_{ac,k}\right)^{\top}}{\sqrt{S}}\right)\to 1$, ${\hat{\bm{y}}}^{i-1}_{na,q}$ and ${\hat{\bm{y}}}^{i-1}_{na,q}$ are very similar. To solve the quadratic complexity, it is necessary to introduce a new operator $\rho$ which avoids the necessity to compute ${{\hat{\bm{y}}}^{i-1}}_{na,q}\times\left({{\hat{\bm{y}}}^{i-1}}_{a,k}\right)^{\top}$ first in practice while retaining the non-negativity. Efficient attention operation [54] is introduced for non-negativity and linear complexity, which employ the softmax operation on $\hat{\bm{y}}^{i-1}_{na}$ in row and the softmax operation on $\hat{\boldmath{y}}^{i-1}_{ac}$ in column.
@@
$\textrm{Attention}=\underbrace{\textrm{softmax}_{2}\left(\hat{\bm{y}}^{i-1}_{na,q}\right)\textrm{softmax}_{1}\left(\hat{\bm{y}}^{i-1}_{ac,k}\right)^{\top}}_{\textrm{non-negative}}\hat{\bm{y}}^{i}_{ac,v}.$ (15)
@@
The process is illustrated in Equation 15. In Equation 15, $\textrm{softmax}_{2}\left(\hat{\bm{y}}^{i-1}_{na,q}\right)\textrm{softmax}_{1}\left(\hat{\bm{y}}^{i-1}_{ac,k}\right)^{\top}$ is employed as the learnable similarity metric, where $10<\textrm{softmax}_{2}\left(\hat{\bm{y}}^{i-1}_{na,q}\right)<1,0<\textrm{softmax}_{1}\left(\hat{\bm{y}}^{i-1}_{ac,k}\right)^{\top}<$ 1, which makes $10<\textrm{softmax}_{2}\left(\hat{\bm{y}}^{i-1}_{na,q}\right)\textrm{softmax}_{1}\left(\hat{\bm{y}}^{i-1}_{ac,k}\right)^{\top}<$ 1. The non-negativity makes $\textrm{softmax}_{2}\left(\hat{\bm{y}}^{i-1}_{na,q}\right)\textrm{softmax}_{1}\left(\hat{\bm{y}}^{i-1}_{ac,k}\right)^{\top}$ can be employed as a similarity metric. If $\textrm{softmax}_{2}\left(\hat{\bm{y}}^{i-1}_{na,q}\right)\textrm{softmax}_{1}\left(\hat{\bm{y}}^{i-1}_{ac,k}\right)^{\top}\to 0$, ${\hat{\bm{y}}}^{i-1}_{na,q}$ and ${\hat{\bm{y}}}^{i-1}_{na,q}$ are near orthogonal. If $\textrm{softmax}_{2}\left(\hat{\bm{y}}^{i-1}_{na,q}\right)\textrm{softmax}_{1}\left(\hat{\bm{y}}^{i-1}_{ac,k}\right)^{\top}\to 1$, ${\hat{\bm{y}}}^{i-1}_{na,q}$ and ${\hat{\bm{y}}}^{i-1}_{na,q}$ are very similar. The metric is implicit because there is no need to compute $\textrm{softmax}_{2}\left(\hat{\bm{y}}^{i-1}_{na,q}\right)\textrm{softmax}_{1}\left(\hat{\bm{y}}^{i-1}_{ac,k}\right)^{\top}$. Since we use softmax operation on $\hat{\bm{y}}^{i-1}_{na,q}$ and $\hat{\bm{y}}^{i-1}_{ac,k}$ separately, $\textrm{softmax}_{1}\left(\hat{\bm{y}}^{i-1}_{ac,k}\right)^{\top}\hat{\bm{y}}^{i}_{ac,v}$ can be computed first during training and testing. The complexity of it is $O(L)$, which is linear with the resolution. The linear complexity makes it easier to employ the global spatial context module for high-resolution image coding.
III-F Linear Complexity Inter-Slice Global Context Module
@@
Because of the global correlations between slices, intra-slice global context module is extended to the inter-slice global context. For a symbol at current slice, the symbol in previous slices is employed at the same position as the approximation of the symbol at current slice, since there are correlations among slices. The correlations among slices or channels are illustrated in Fig. 2 and Fig. 8. This approximation makes the anchor part and non-anchor part benefit from more contexts. The process of $i$-th slice is taken as an example. Same as linear complexity intra-slice global context module, attention mechanism is employed to measure the similarity. To make the inter-slice global context capturing more efficient, the softmax operation in vanilla attention is divided into two independent softmax operations as discussed in Section III-E. The overall process is
$\hat{\bm{y}}^{i}_{attn}=\underbrace{\textrm{softmax}_{2}\left(\hat{\bm{y}}^{<i}_{q}\right)\textrm{softmax}_{1}\left(\hat{\bm{y}}^{<i}_{k}\right)^{\top}}_{\textrm{non-negative}}\hat{\bm{y}}^{<i}_{v},$ (16)
$\hat{\bm{y}}^{i}_{conv}=\textrm{conv}_{K\times K}(\hat{\bm{y}}^{i}_{attn})+\hat{\bm{y}}^{i}_{a,v},$ (17)
${\bm{\Phi}}^{i}_{gc,inter}=\textrm{DepthRB}(\hat{\bm{y}}^{i}_{conv}),$ (18)
where $\hat{\bm{y}}^{<i}_{q},\hat{\bm{y}}^{<i}_{k},\hat{\bm{y}}^{<i}_{v}=\textrm{Embedding}(\hat{\bm{y}}^{<i})$, Embedding is the embedding layer. Embedding layer consists of a $1\times 1$ convolutional layer and a $3\times 3$ depth-wise convolutional layer. The $3\times 3$ depth-wise convolutional layer is employed for learnable position embedding. DepthRB is the depth-wise residual bottleneck [36] and is employed to enhance the non-linearity. Since the $\textrm{softmax}_{1}\left(\hat{\bm{y}}^{<i}_{k}\right)^{\top}\hat{\bm{y}}^{<i}_{v}$ is computed during training and testing, the similarity metric is implicit and the overall complexity is $O(L)$, which is linear with the resolution.
Methods Kodak [55] Tecnick [56] CLIC Professional Valid [22]
@@
VTM-17.0 Intra [5] $0.00$ $0.00$ $0.00$ $0.00$ $0.00$ $0.00$
Cheng’20 (CVPR’20) [24] $+5.58$ $-44.21$ $+7.57$ $-39.61$ $+11.71$ $-41.29$
Minnen’20 (ICIP’20) [23] $+3.23$- $-0.88$- -- --
Qian’21 (ICLR’21) [26] $+10.05$ $-39.53$ $+7.52$- $+0.02$-
Xie’21 (ACMMM’21) [28] $+1.55$ $-43.39$ $+3.21$- $+0.99$-
Guo’22 (TCSVT’22) [14] $-4.45$ $-45.23$- -- -- --
LBHIC (TCSVT’22) [6] $-4.56$ $-50.54$- -- -- --
Entroformer (ICLR’22) [41] $+4.73$ $-42.64$ $+2.31$- $-1.04$-
SwinT-Charm (ICLR’22) [38] $-1.73$- -- -- -- --
NeuralSyntax (CVPR’22) [57] $+8.97$ $-39.56$- -- $+5.64$ $-38.92$
McQuic (CVPR’22) [58] $-1.57$ $-47.94$- -- $+6.82$ $-40.17$
STF (CVPR’22) [25] $-2.48$ $-47.72$ $-2.75$- $+0.42$-
WACNN (CVPR’22) [25] $-2.95$ $-47.71$ $-5.09$- $+0.04$-
ELIC (CVPR’22) [29] $-5.95$ $-44.60$ $-9.14$- $-3.45$-
Contextformer (ECCV’22) [34] $-5.77$ $-46.12$ $-9.05$ $-42.29$- --
Pan’22 (ECCV’22) [59] $+7.56$ $-36.20$ $+3.97$- -- --
NVTC (CVPR’23) [60] $-1.04$- -- -- $-3.61$-
LIC-TCM (CVPR’23) [27] $-10.14$ $-48.94$ $-11.47$- $-8.04$-
MLIC (ACMMM’23) [30] $-8.05$ $-49.13$ $-12.73$ $-47.26$ $-8.79$ $-45.79$
${}^{+}$ (ACMMM’23) [30] $-11.39$ $-52.75$ $-16.38$ $-53.54$ $-12.56$ $-48.75$
@@
Figure 11: Reconstructions of our proposed ${}^{++}$, recent learned image compression models [29, 25, 28, 24] and VTM-17.0 Intra [5]. “[MSE]” denotes the model is optimized for MSE and “[MS-SSIM]” denotes the model is optimized for MS-SSIM [61]. Please zoom in for better view.
IV Experiments
IV-A Implementation Details
IV-A1 Training Dataset
@@
Our training datasets contains $10^{5}$ images [^1] https://github.com/JiangWeibeta/MLIC/blob/main/train_list.txt with resolutions exceeding $512\times 512$. These images are selected from ImageNet [62], COCO 2017 [63], DIV2K [64], and Flickr2K [65]. To address the existing compression artifacts in JPEG images, we follow the approach of Ballé et al [17] by further down-sampling the JPEG images using a randomized factor. This downsampling process ensures that the minimum height or width of the images falls within the range of 512 to 584 pixels, effectively reducing the compression artifacts.
IV-A2 Training Strategy
@@
${}^{++}$ is built on Pytorch [66] and CompressAI [67]. Following the settings of CompressAI [67], we set $\lambda\in\{18,35,67,130,250,483\}\times 10^{-4}$ for MSE and set $\lambda\in\{2.4,4.58,8.73,16.64,31.73,60.5\}$ for Multi-Scale Structural Similarity (MS-SSIM) [61]. The batch size is set to $32$ and models are trained on a single Tesla A100 GPU. Each model is trained with an Adam optimizer with $\beta_{1}=0.9,\beta_{2}=0.999$. We train each model for 2M steps. The learning rate starts at $10^{-4}$ and drops to $3\times 10^{-5}$ at 1.5M steps, drops to $10^{-5}$ at 1.8M steps, and drops to $3\times 10^{-6}$ at 1.9M steps, drops to $10^{-6}$ at 1.95M steps. During training, we random crop images to $256\times 256$ patches during the first $1.2$M steps. To further exploit the effectiveness of global context modules, images are cropped to $512\times 512$ patches during the rest steps. Large patches are beneficial for learning global references. The latent representation can be sparse due to checkerboard partition. A large latent representation makes it more difficult for the model to capture the global contexts and improves the generalization ability of the model across different resolutions.
IV-B Benchmarks and Metrics

### 20 Feb 2024 v9

@@ 3.5 Linear Complexity Intra-Slice Global Context Module
One drawback of vanilla approach is its quadratic complexity. In Equation 12, the softmax operation specifies the order of tensor calculation. The complexity of ${{\hat{\boldsymbol{y}}}^{i-1}}_{na,q}\times\left({{\hat{\boldsymbol{y}}}^{i-1}}_{a,k}\right)^{\top}$ is $O(L^{2})$. The quadratic complexity leads to huge GPU memory consumption, longer encoding and decoding time as illustrated in Fig. 12, which makes it hard to employ the vanilla approach for high-resolution image coding. In Equation 12, if $\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}\times\hat{\boldsymbol{y}}^{i}_{ac,v}$ is computed first, the overall complexity becomes $O(L)$, which is linear with the resolution. Equation 12 works because $10<\textrm{softmax}\left(\frac{{{\hat{\boldsymbol{y}}}^{i-1}}_{na,q}\times\left({{\hat{\boldsymbol{y}}}^{i-1}}_{ac,k}\right)^{\top}}{\sqrt{S}}\right)<$ 1. The non-negativity makes it can be treated as a learnable similarity metric. If $\textrm{softmax}\left(\frac{{{\hat{\boldsymbol{y}}}^{i-1}}_{na,q}\times\left({{\hat{\boldsymbol{y}}}^{i-1}}_{ac,k}\right)^{\top}}{\sqrt{S}}\right)\to 0$, ${\hat{\boldsymbol{y}}}^{i-1}_{na,q}$ and ${\hat{\boldsymbol{y}}}^{i-1}_{na,q}$ are near orthogonal. If $\textrm{softmax}\left(\frac{{{\hat{\boldsymbol{y}}}^{i-1}}_{na,q}\times\left({{\hat{\boldsymbol{y}}}^{i-1}}_{ac,k}\right)^{\top}}{\sqrt{S}}\right)\to 1$, ${\hat{\boldsymbol{y}}}^{i-1}_{na,q}$ and ${\hat{\boldsymbol{y}}}^{i-1}_{na,q}$ are very similar. To solve the quadratic complexity, it is necessary to introduce a new operator $\rho$ which avoids the necessity to compute ${{\hat{\boldsymbol{y}}}^{i-1}}_{na,q}\times\left({{\hat{\boldsymbol{y}}}^{i-1}}_{a,k}\right)^{\top}$ first in practice while retaining the non-negativity. Efficient attention operation (Shen et al., 2021) is introduced for non-negativity and linear complexity, which employ the softmax operation on $\hat{\boldsymbol{y}}^{i-1}_{na}$ in row and the softmax operation on $\hat{\boldmath{y}}^{i-1}_{ac}$ in column.
@@
$\textrm{Attention}=\underbrace{\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i-1}_{na,q}\right)\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}}_{\textrm{non-negative}}\hat{\boldsymbol{y}}^{i}_{ac,v}.$ (15)
@@
The process is illustrated in Equation 15. In Equation 15, $\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i-1}_{na,q}\right)\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}$ is employed as the learnable similarity metric, where $10<\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i-1}_{na,q}\right)<1,0<\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}<$ 1, which makes $10<\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i-1}_{na,q}\right)\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}<$ 1. The non-negativity makes $\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i-1}_{na,q}\right)\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}$ can be employed as a similarity metric. If $\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i-1}_{na,q}\right)\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}\to 0$, ${\hat{\boldsymbol{y}}}^{i-1}_{na,q}$ and ${\hat{\boldsymbol{y}}}^{i-1}_{na,q}$ are near orthogonal. If $\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i-1}_{na,q}\right)\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}\to 1$, ${\hat{\boldsymbol{y}}}^{i-1}_{na,q}$ and ${\hat{\boldsymbol{y}}}^{i-1}_{na,q}$ are very similar. The metric is implicit because there is no need to compute $\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i-1}_{na,q}\right)\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}$. Since we use softmax operation on $\hat{\boldsymbol{y}}^{i-1}_{na,q}$ and $\hat{\boldsymbol{y}}^{i-1}_{ac,k}$ separately, $\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}\hat{\boldsymbol{y}}^{i}_{ac,v}$ can be computed first during training and testing. The complexity of it is $O(L)$, which is linear with the resolution. The linear complexity makes it easier to employ the global spatial context module for high-resolution image coding.
3.6 Linear Complexity Inter-Slice Global Context Module
@@
Because of the global correlations between slices, intra-slice global context module is extended to the inter-slice global context. For a symbol at current slice, the symbol in previous slices is employed at the same position as the approximation of the symbol at current slice, since there are correlations among slices. The correlations among slices or channels are illustrated in Fig. 2 and Fig. 8. This approximation makes the anchor part and non-anchor part benefit from more contexts. The process of $i$-th slice is taken as an example. Same as linear complexity intra-slice global context module, attention mechanism is employed to measure the similarity. To make the inter-slice global context capturing more efficient, the softmax operation in vanilla attention is divided into two independent softmax operations as discussed in Section 3.5. The overall process is
$\hat{\boldsymbol{y}}^{i}_{attn}=\underbrace{\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{<i}_{q}\right)\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{<i}_{k}\right)^{\top}}_{\textrm{non-negative}}\hat{\boldsymbol{y}}^{<i}_{v},$ (16)
$\hat{\boldsymbol{y}}^{i}_{conv}=\textrm{conv}_{K\times K}(\hat{\boldsymbol{y}}^{i}_{attn}),$ (17)
${\boldsymbol{\Phi}}^{i}_{gc,inter}=\textrm{DepthRB}(\hat{\boldsymbol{y}}^{i}_{conv}),$ (18)
where $\hat{\boldsymbol{y}}^{<i}_{q},\hat{\boldsymbol{y}}^{<i}_{k},\hat{\boldsymbol{y}}^{<i}_{v}=\textrm{Embedding}(\hat{\boldsymbol{y}}^{<i})$, Embedding is the embedding layer. Embedding layer consists of a $1\times 1$ convolutional layer and a $3\times 3$ depth-wise convolutional layer. The $3\times 3$ depth-wise convolutional layer is employed for learnable position embedding. DepthRB is the depth-wise residual bottleneck (Jiang et al., 2023a) and is employed to enhance the non-linearity. Since the $\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{<i}_{k}\right)^{\top}\hat{\boldsymbol{y}}^{<i}_{v}$ is computed during training and testing, the similarity metric is implicit and the overall complexity is $O(L)$, which is linear with the resolution.
Methods Kodak (Kodak, 1993) Tecnick (Asuni & Giachetti, 2014) CLIC Professional Valid (Toderici et al., 2020)
@@
VTM-17.0 Intra (Bross et al., 2021) $0.00$ $0.00$ $0.00$ $0.00$ $0.00$ $0.00$
Cheng’20 (CVPR’20) (Cheng et al., 2020) $+5.58$ $-44.21$ $+7.57$ $-39.61$ $+11.71$ $-41.29$
Minnen’20 (ICIP’20) (Minnen & Singh, 2020) $+3.23$- $-0.88$- -- --
Qian’21 (ICLR’21) (Qian et al., 2020) $+10.05$ $-39.53$ $+7.52$- $+0.02$-
Xie’21 (ACMMM’21) (Xie et al., 2021) $+1.55$ $-43.39$ $+3.21$- $+0.99$-
Guo’22 (TCSVT’22) (Guo et al., 2022) $-4.45$ $-45.23$- -- -- --
LBHIC (TCSVT’22) (Wu et al., 2022) $-4.56$ $-50.54$- -- -- --
Entroformer (ICLR’22) (Qian et al., 2022) $+4.73$ $-42.64$ $+2.31$- $-1.04$-
SwinT-Charm (ICLR’22) (Zhu et al., 2022b) $-1.73$- -- -- -- --
NeuralSyntax (CVPR’22) (Wang et al., 2022) $+8.97$ $-39.56$- -- $+5.64$ $-38.92$
McQuic (CVPR’22) (Zhu et al., 2022a) $-1.57$ $-47.94$- -- $+6.82$ $-40.17$
STF (CVPR’22) (Zou et al., 2022) $-2.48$ $-47.72$ $-2.75$- $+0.42$-
WACNN (CVPR’22) (Zou et al., 2022) $-2.95$ $-47.71$ $-5.09$- $+0.04$-
ELIC (CVPR’22) (He et al., 2022) $-5.95$ $-44.60$ $-9.14$- $-3.45$-
Contextformer (ECCV’22) (Koyuncu et al., 2022) $-5.77$ $-46.12$ $-9.05$ $-42.29$- --
Pan’22 (ECCV’22) (Pan et al., 2022) $+7.56$ $-36.20$ $+3.97$- -- --
NVTC (CVPR’23) (Feng et al., 2023) $-1.04$- -- -- $-3.61$-
LIC-TCM (CVPR’23) (Liu et al., 2023) $-10.14$ $-48.94$ $-11.47$- $-8.04$-
MLIC (ACMMM’23) (Jiang et al., 2023b) $-8.05$ $-49.13$ $-12.73$ $-47.26$ $-8.79$ $-45.79$
${}^{+}$ (ACMMM’23) (Jiang et al., 2023b) $-11.39$ $-52.75$ $-16.38$ $-53.54$ $-12.56$ $-48.75$
@@
Figure 11: Reconstructions of our proposed ${}^{++}$, recent learned image compression models (He et al., 2022; Zou et al., 2022; Xie et al., 2021; Cheng et al., 2020) and VTM-17.0 Intra (Bross et al., 2021). “[MSE]” denotes the model is optimized for MSE and “[MS-SSIM]” denotes the model is optimized for MS-SSIM (Wang et al., 2003). Please zoom in for better view.
4 Experiments
4.1 Implementation Details
4.1.1 Training Dataset
@@
Our training datasets contains $10^{5}$ images [^1] https://github.com/JiangWeibeta/MLIC/blob/main/train_list.txt with resolutions exceeding $512\times 512$. These images are selected from ImageNet (Deng et al., 2009), COCO 2017 (Lin et al., 2014), DIV2K (Agustsson & Timofte, 2017), and Flickr2K (Lim et al., 2017). To address the existing compression artifacts in JPEG images, we follow the approach of Ballé et al (Ballé et al., 2018) by further down-sampling the JPEG images using a randomized factor. This downsampling process ensures that the minimum height or width of the images falls within the range of 512 to 584 pixels, effectively reducing the compression artifacts.
4.1.2 Training Strategy
@@
${}^{++}$ is built on Pytorch (Paszke et al., 2019) and CompressAI (Bégaint et al., 2020). Following the settings of CompressAI (Bégaint et al., 2020), we set $\lambda\in\{18,35,67,130,250,483\}\times 10^{-4}$ for MSE and set $\lambda\in\{2.4,4.58,8.73,16.64,31.73,60.5\}$ for Multi-Scale Structural Similarity (MS-SSIM) (Wang et al., 2003). The batch size is set to $32$ and models are trained on a single Tesla A100 GPU. Each model is trained with an Adam optimizer with $\beta_{1}=0.9,\beta_{2}=0.999$. We train each model for 2M steps. The learning rate starts at $10^{-4}$ and drops to $3\times 10^{-5}$ at 1.5M steps, drops to $10^{-5}$ at 1.8M steps, and drops to $3\times 10^{-6}$ at 1.9M steps, drops to $10^{-6}$ at 1.95M steps. During training, we random crop images to $256\times 256$ patches during the first $1.2$M steps. To further exploit the effectiveness of global context modules, images are cropped to $512\times 512$ patches during the rest steps. Large patches are beneficial for learning global references. The latent representation can be sparse due to checkerboard partition. A large latent representation makes it more difficult for the model to capture the global contexts and improves the generalization ability of the model across different resolutions.
4.2 Benchmarks and Metrics